OpenStudy (anonymous):

3(x-2)^3/4=24 may be two answers to this one

OpenStudy (anonymous):

are you repeating yourself? i thought we did this one last night

OpenStudy (anonymous):

it was a similar one but i can't find this exact one

OpenStudy (anonymous):

it was this one (x+1)^3/2-2=25 and the answer was 8

OpenStudy (anonymous):

\[(x-2)^{\frac{3}{4}}=8\] \[x-2=8^{\frac{4}{3}}=16\] \[x=18\]

OpenStudy (anonymous):

\[x-2=-16\] \[x=-14\]